stion 13 of 14 aping learning For a reaction, stan
Answer #1

G = G° + RT*lnKc

G° = -32.8

RT = 8.314*298

Kp = CH4^2 / (C2H6)(H2)

Kp = 0.7^2 / (0.45*0.45) = 2.41975

dn = 2 -1-1 = 0

Kp = Kc*(RT)^dn

Kp = Kc

then

G = G° + RT*lnKc

G= -32800 + 8.314*298*ln(2.41975) = -30610.6 J

G = -30.610 kJ/mol